# Gravity as a topological gauge theory

### v1

@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can be decomposed as :
where $A=\dfrac{1}{2}A^{ab}J_{ab}=\dfrac{1}{2}A^{ab}_{\mu}dx^{\mu}\otimes J_{ab}$ corresponds to the spin connection and $\beta=\beta^{a}e_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes e_{a}$ corresponds to the tetrad or soldering form used to define the metric of GR $\tilde{g}=\eta(\beta,\beta)$. $\alpha=\alpha^{a}e_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes e_{a}$ can be interpreted as a secondary tetrad. Finally $\lambda=\lambda_{\mu}dx^{\mu}$ is the $\mathbb{R}$-valued (dilations) part of the connection. Also : $\bar{\alpha}=\alpha^{a}\bar{e}_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$ and $\bar{\beta}=\beta^{a}\bar{e}_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$. The expression of the curvature $\bar{\Omega}=d\varpi+\dfrac{1}{2}[\varpi,\varpi]$ of $\varpi$ is :
@@
The Bianchi identities are given by $D\bar{\Omega}=d\bar{\Omega}+\varpi\wedge\bar{\Omega}-\bar{\Omega}\wedge\varpi=0$ which splits into the identities:
@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
$S_{G}[\varpi]$ $=e\mathdutchcal{E}+r\mathdutchcal{P}=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})\Bigr{)}$ (1)
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}F^{ab}\wedge F^{cd}-\dfrac{r}{8\pi^{2}}Tr(\bar{\Omega}\wedge\bar{\Omega})\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi^{cd}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\phi^{ab}\wedge\phi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\underbrace{\phi^{ab}\wedge\phi_{ab}}_{=0}-\dfrac{4}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2f\wedge f\bigr{)}\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{4}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon_{abcd}\varepsilon^{r_{1}s_{1}cs_{2}}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;r_{1}}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;s_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{4}}\beta^{\,\mu_{4}}_{~{}\,\,s_{2}}d^{4}x$
@@
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{4}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon^{r_{1}s_{1}r_{2}s_{2}}R^{ab}_{\;\;\;r_{1}s_{1}}\beta^{\,c}_{~{}\,\,\mu_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{2}}\beta^{-1\,\mu_{1}}_{~{}~{}\,\,\;\;\;r_{2}}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;s_{2}}\eta_{ca}\eta_{db}d^{4}x$
$\qquad\qquad\quad\quad-\dfrac{4}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}.$
@@
$S_{G}[\varpi]$ $=e\mathdutchcal{E}+r\mathdutchcal{P}=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\mathscr{M})\Bigr{)}=\int_{\mathscr{M}}\Bigl{(}e\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{so}(3,1)}}{2\pi})+rP(\bar{\Omega})\Bigr{)}$ (2)
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi^{cd}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\xi^{ab}\wedge\xi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\underbrace{\xi^{ab}\wedge\xi_{ab}}_{=0}-\dfrac{2k}{\ell\ell^{\prime}}\eta_{ab}\Theta^{a}\wedge\Theta^{b}-2f^{2}\bigr{)}\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi^{cd}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\xi^{ab}\wedge\xi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}-\dfrac{2k}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-\dfrac{4k}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{8k}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;b}d^{4}x+\dfrac{24k^{2}}{\ell^{2}\ell^{\prime 2}}\sqrt{|\tilde{g}|}d^{4}x\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon^{rs}_{~{}~{}ab}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;r}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;s}d^{4}x$
$\qquad\qquad\quad\quad-\dfrac{2k}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-\dfrac{4k}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}$
@@
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}$ (3)
which corresponds to the action (4) for $G=ISO(3,1)\Leftrightarrow k=0$, up to the $\lambda$ kinetic term. However, ignoring the boundary terms leaves the equations of motion unchanged if we keep the same matter Lagrangian as in the preceding example 2. While for $k\neq 0$ a direct comparison between $S^{\prime}$ (3) and $S_{G}$ (2) amounts to identifying the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{1}\\ \dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{5}\\ \dfrac{ek^{2}}{8(\pi\ell\ell^{\prime})^{2}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{2\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{3}=-2\alpha_{5}\\ \alpha_{1}=-\dfrac{2ek}{r\ell\ell^{\prime}}\alpha_{3}\\ r=\dfrac{2\pi^{2}\ell\ell^{\prime}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (4)
@@
$\begin{cases}\alpha_{1}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{2\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{2k}{\ell\ell^{\prime}}\alpha_{3}=-\alpha_{5}\\ \alpha_{1}=\dfrac{e}{r}\alpha_{2}\\ r=\dfrac{4\pi^{2}\ell\ell^{\prime}}{k}\alpha_{2}\end{cases}$ (5)
@@
$S_{G}[\varpi]=\int_{\mathscr{M}}$ $\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell\ell^{\prime}}(\dfrac{e}{2}R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+rR^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Holst}+\overbrace{\dfrac{2ek^{2}}{(4\pi\ell\ell^{\prime})^{2}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad+\,\overbrace{\dfrac{r}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}$

### v3

@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can accordingly be decomposed as:
where $A=\dfrac{1}{2}A^{ab}J_{ab}=\dfrac{1}{2}A^{ab}_{\mu}dx^{\mu}\otimes J_{ab}$ corresponds to the spin connection and $\beta=\beta^{a}e_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes e_{a}$ corresponds to the tetrad or soldering form used to define the metric of the tetrad formalism $\tilde{g}=\eta(\beta,\beta)$. $\alpha=\alpha^{a}e_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes e_{a}$ can be interpreted as a secondary tetrad. Finally $\lambda=\lambda_{\mu}dx^{\mu}$ is the $\mathbb{R}$-valued (dilations) part of the connection while $\ell$ and $\ell^{\prime}$ are non null real numbers. Also : $\bar{\alpha}=\alpha^{a}\bar{e}_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$ and $\bar{\beta}=\beta^{a}\bar{e}_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$.
The curvature $\bar{\Omega}=d\varpi+\dfrac{1}{2}[\varpi,\varpi]$ of $\varpi$ then reads:
@@
The Bianchi identities are given by $D\bar{\Omega}=d\bar{\Omega}+\varpi\wedge\bar{\Omega}-\bar{\Omega}\wedge\varpi=0$ which can be split into the identities:
@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
$S_{G}[\varpi]$ $=e\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+y\mathdutchcal{P}(F)=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})+yP(F)\Bigr{)}$ (1)
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}F^{ab}\wedge F^{cd}-\dfrac{r}{8\pi^{2}}Tr(\bar{\Omega}\wedge\bar{\Omega})-\dfrac{y}{8\pi^{2}}Tr(F\wedge F)\Bigr{)}$
$=\int_{\mathscr{M}}\biggl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi^{cd}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\phi^{ab}\wedge\phi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\underbrace{\phi^{ab}\wedge\phi_{ab}}_{=0}\bigr{)}-\dfrac{4r}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2rf\wedge f\Bigr{)}\biggr{)}$
$=\int_{\mathscr{M}}\biggl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{4}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon_{abcd}\varepsilon^{r_{1}s_{1}cs_{2}}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;r_{1}}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;s_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{4}}\beta^{\,\mu_{4}}_{~{}\,\,s_{2}}d^{4}x$
@@
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{4}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon^{r_{1}s_{1}r_{2}s_{2}}R^{ab}_{\;\;\;r_{1}s_{1}}\beta^{\,c}_{~{}\,\,\mu_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{2}}\beta^{-1\,\mu_{1}}_{~{}~{}\,\,\;\;\;r_{2}}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;s_{2}}\eta_{ca}\eta_{db}d^{4}x\bigr{)}$
$\qquad\qquad\quad\quad-\dfrac{4r}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2rf\wedge f\Bigr{)}\biggr{)}.$
@@
$S_{G}[\varpi]$ $=e\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+y\mathdutchcal{P}(F)=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})+yP(F)\Bigr{)}=\int_{\mathscr{M}}\Bigl{(}e\operatorname{Pf}(\dfrac{F}{2\pi})+rP(\bar{\Omega})+yP(F)\Bigr{)}$ (2)
$=\int_{\mathscr{M}}\biggl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi^{cd}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\xi^{ab}\wedge\xi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\underbrace{\xi^{ab}\wedge\xi_{ab}}_{=0}\bigr{)}-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}\Theta^{a}\wedge\Theta^{b}-2rf^{2}\Bigr{)}\biggr{)}$
$=\int_{\mathscr{M}}\biggl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi^{cd}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\xi^{ab}\wedge\xi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}\bigr{)}-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-\dfrac{4rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}-2rd\lambda\wedge d\lambda\Bigr{)}\biggr{)}$
$=\int_{\mathscr{M}}\biggl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{8k}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;b}d^{4}x+\dfrac{24k^{2}}{\ell^{2}\ell^{\prime 2}}\sqrt{|\tilde{g}|}d^{4}x\bigr{)}$
$\qquad\quad+\dfrac{1}{8\pi^{2}}\Bigl{(}(r+y)\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon^{rs}_{~{}~{}ab}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;r}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;s}d^{4}x\bigr{)}$
$\qquad\qquad\quad\quad-\dfrac{2rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-\dfrac{4rk}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}-2rd\lambda\wedge d\lambda\Bigr{)}\biggr{)}$
@@
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{1}{8\pi^{2}}\bigl{(}(r+y)R^{ab}\wedge R_{ab}-2rd\lambda\wedge d\lambda\bigr{)}\Bigr{)}.$ (3)
This corresponds to the action (5) for $G=ISO(3,1)\Leftrightarrow k=0$, up to the $\lambda$ kinetic term. However, ignoring the boundary terms leaves the equations of motion unchanged if we keep the same matter Lagrangian as in the preceding example 3.
While for $k\neq 0$ a direct comparison between $S^{\prime}$ (4) and $S_{G}$ (2) amounts to identifying the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{1}\\ \dfrac{(r+y)k}{2\pi^{2}\ell\ell^{\prime}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r+y}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{5}\\ \dfrac{ek^{2}}{8(\pi\ell\ell^{\prime})^{2}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{2\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{3}=-2(1+\dfrac{y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{2ek}{(r+y)\ell\ell^{\prime}}\alpha_{3}\\ r+y=\dfrac{2\pi^{2}\ell\ell^{\prime}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (4)
@@
$\begin{cases}\alpha_{1}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{2\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{k(r+2y)}{4\pi^{2}\ell\ell^{\prime}}=\dfrac{2k(r+2y)}{(r+y)\ell\ell^{\prime}}\alpha_{3}=-(1+\dfrac{2y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{e}{r+2y}\alpha_{2}\end{cases}$ (5)
@@
$S_{G}[\varpi]=\int_{\mathscr{M}}$ $\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell\ell^{\prime}}\bigl{(}\dfrac{e}{2}R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(r+2y)R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Holst}+\overbrace{\dfrac{2ek^{2}}{(4\pi\ell\ell^{\prime})^{2}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad+\,\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}$